Show work please. consider the functions f(x)=4x-16 and g(x)= x/4 +4.
Find f(g(x)) and g(f(x)) and determine whether the functions f and g are inverses of each other.
thanks so much!

Answer by josgarithmetic(39618)

4x-16, and x/4+4

This substitution,
4(x/4+4)-16=x+16-16
=x

The other substitution,
(4x-16)/4+4=(x-4)+4
=x-4+4
=x